### v1

# ChatWise: AI-Powered Engaging Conversations for Enhancing Senior Cognitive Wellbeing

Abstract
Cognitive health in older adults presents a growing challenge. While conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overburdened. AI-based methods have shown promise in providing conversational support, yet existing work is limited to implicit strategy while lacking multi-turn support tailored to seniors. We improve prior art with an LLM-driven chatbot named ChatWise  for older adults. It follows dual-level conversation reasoning at the inference phase to provide engaging companionship. ChatWise thrives in long-turn conversations, in contrast to conventional LLMs that primarily excel in short-turn exchanges. Grounded experiments show that ChatWise significantly enhances simulated users’ cognitive and emotional status, including those with Mild Cognitive Impairment.
ChatWise: AI-Powered Engaging Conversations
for Enhancing Senior Cognitive Wellbeing
Zhengbang Yang, Zhuangdi Zhu George Mason University, Fairfax, VA, USA {zyang30,zzhu24}@gmu.edu
1 Introduction
[…]
In this work, we aim to draw on the advance of AI to provide older adults with engaging conversational support that improves their cognitive function and reduces social loneliness, serving as accessible alternatives to human companions. Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), revealing that conversation behavior can have a measurable influence on interlocutors, which inspired us to develop principled yet efficient methods that transform clinical insights into strategic dialogue design leveraging the advanced reasoning ability of LLMs.
In response, we propose an LLM-driven chatbot named ChatWise. Unlike traditional chatbot driven with implicit conversational goals, ChatWise employs dual-layer conversation generation, which first derives categorized macro-level information, including user emotion states to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions. Figure 1 overviews its development design. Our work provides multifold contributions:
  • ChatWise is a tuning-free yet principled framework for developing AI chatbots with enhanced multi-turn, daily-themed conversational support. Our dual-level conversation generation design integrates clinical insights into LLM reasoning and can be readily applied to various LLMs.
  • ChatWise follows grounded development and evaluation using digital twins (Hong et al., 2024), which are simulated users constructed by tuning LLMs on de-identified, real-world dialogues between seniors and professional caregivers of a clinical trial (I-CONECT, 2024). Interactions with digital twins demonstrated that ChatWise significantly enhances users’ engagement and cognitive status with multi-turn interactions, especially for users with Mild Cognitive Impairment (MCI) (Figure 2).
  • Our comparative studies demonstrated that integrating macro-level strategies into conversation generation is the key contributor to enhancing user engagement. Its gain on cognitive metrics remains consistent across different backbone LLMs. ChatWise’s advantages become more pronounced in longer conversations, in contrast to conventional LLMs that primarily excel in short-turn question-answer exchanges.
  • Our dialogue analysis revealed key empirical insights, including the pattern of predominant interaction strategies across users. While transient user emotional fluctuations were less tractable, multi-turn conversations with ChatWise have shown to notably improve user emotions over time. This highlights the value of long-term, principled conversational support in enhancing seniors’ cognitive and emotional well-being. We hope these findings can inspire future research, including RL-based LLM optimization.
Figure 1: Overview of ChatWise Development.
Figure 2: ChatWise notably improves multiple user cognitive metrics through 10-turn dialogues (Sec 4.2).
2 Related Work
AI-powered Chatbots for Seniors: Social interaction plays a crucial role in the cognitive health of seniors  (Patil and Braun, 2024; Fiori and Jager, 2012). AI-powered chatbots have emerged as a promising alternative  (Rodríguez-Martínez et al., 2023; Owan et al., 2023). Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance  (Yang et al., 2024; Liu et al., 2021b; sen, 2024; Ryu et al., 2020; Hong et al., 2024). Unlike prior work, our approach prioritizes user conversational engagement, which has been empirically shown to enhance both cognitive and emotional status.
Conversational Strategies: Liu et al. (2021a) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support. Yuan et al. (2023) examined the causal relationships between dialogue acts (DAs) and participants’ emotional states in a clinical trial (I-CONECT, 2024), emphasizing the impact of strategic interventions in tele-mediated dialogues. Seo et al. (2021) identified key strategies for improving child patient-provider communication through semi-structured interviews. Few works have systematically integrated these strategies into the reasoning flow of AI chatbots. Our approach fills this gap by embedding structured conversational strategies for automatically enhancing user engagement.
Multi-turn chatbot exploration: Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024). While some pioneering research has explored multi-turn optimization through Reinforcement Learning (RL) approaches for LLMs (Verma et al., 2022; Zhou et al., 2024; Abdulhai et al., 2023; Gao et al., 2025), these methods were not specifically designed for supporting senior dialogue engagement. Our approach is tuning-free yet compatible with RL-driven methods, enabling future extensions through hierarchical RL to further enhance multi-turn conversation engagement.
3 ChatWise Design
ChatWise employs a dual-level framework with a strategy provider $\pi_{s}$ that induces macro-level actions, and an utterance generator $\pi_{u}$ for deriving utterances based on suggested strategies (Figure 1).
Strategy Candidates: To construct a strategy candidate set, we extract Dialogue Acts (DAs) from real telehealth clinical trials Yuan et al. (2023), where DAs serve as atomic communicative units that convey distinct conversational intentions Searle (1976). We further integrate these with strategies from prior emotional support dataset (Liu et al., 2021a) to form a comprehensive set. We refine each strategy with definitions and examples as in-context prompts to enhance $\pi_{s}$ reasoning.
From Strategy to Utterance: $\pi_{s}$ reasons over the dialogue history $\tau$ to capture user’s emotion $e$ and derive one to two strategies, $\{s_{i}\}_{i\leq 2}\sim\pi_{s}(\tau,e)$. Our rationale follows counseling study (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g. initiating new topics via an open question, or backward, e.g. responding with acknowledgment. $\pi_{s}$ is prompted to flexibly employ one or both as needed. Selected strategies, user’s current emotion, and conversation history serve as inputs for $\pi_{u}$ to generate utterances. To ensure natural flow, $\pi_{u}$ will first improvise a few rounds before following suggested strategies. Drawing on clinical guidance, $\pi_{u}$ encourages users to choose topics rather than imposing them.
4 Empirical Evaluation
Below we summarize experiment configurations, metrics, and main findings of ChatWise. More experimental details are available in Appendix A.1 and A.2.
4.1 Conversation Generation
For controlled and responsive development of ChatWise, we employed 9 digital twins provided by  Hong et al. (2024) as simulated users with different personas, which are fine-tuned LLMs to approximate senior individuals with MCI symptoms using real, de-identified dialogue data from a clinical trial  I-CONECT (2024). We collected 20 trajectories of conversations between ChatWise and each digital twin, where a conversation contains 10 turns, each turn with a two-way utterance exchange between ChatWise and digital twin (user).
4.2 Engagement Evaluation
User verbosity: We primarily measure user engagement through verbosity, i.e. the user-to-chatbot token ratio, defined as $\frac{\text{tokens}_{user}}{\text{tokens}_{chatbot}}$, as a clinical study (Yu et al., 2021) suggested reducing the moderator talkativeness while encouraging participant (user) expression.
Cognitive Metrics: See et al. (2019) defines 9 aspects for evaluating conversation quality, from which we select 5 focused on engagement assessment: Engagingness, Interestingness, Listening, Fluency, and Making Sense. While Interestingness assesses overall dialogue quality, the others primarily evaluate user interactions.
We compute the win rate (Rafailov et al., 2023) by comparing dialogues generated with and without ChatWise, using the same pretrained LLM (GPT-4o) as the utterance generator and identical system prompts, except for input from $\pi_{s}$ to $\pi_{u}$. Dialogue pairs are randomly matched within the same digital twin for evaluation.
Strategy Provider Verbosity $\uparrow$ Engagingness WR $\uparrow$ Interestingness WR $\uparrow$ Listening WR $\uparrow$ Fluency WR $\uparrow$ Making sense WR $\uparrow$
w/o ChatWise 0.7398 0.3389 0.3833 0.3593 0.4111 0.4056
GPT-4o 0.8635 0.6833 0.6222 0.6222 0.5778 0.5833
o3-mini 0.8643 0.6778 0.6222 0.6778 0.6222 0.6167
Llama 3.1-405B 0.8083 0.6222 0.6056 0.6222 0.5667 0.5833
Table 1: Performance on 10-turn dialogue across different strategy providers. W/o ChatWise is the baseline without applying any strategy, whose win rates (WRs) are averaged across strategy providers. (Appendix A.3)
4.3 ChatWise Performance and Analysis
We evaluated ChatWise using different LLM backbones as strategy providers, including GPT-4o, o3-mini, and Llama3.1-405B. As shown in Table 1, ChatWise consistently outperformed non-strategic dialogues across all evaluation metrics.
ChatWise robustly enhances user engagement: Regardless of the capability of pretrained LLMs, our design consistently improves user engagement and dialogue quality with a large margin across all tested LLMs as the strategy provider.
Reasoning ability can transfer to improve engagement: The best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
4.4 Accumulated Multi-Turn Analysis
We analyzed ChatWise’s performance over increasing conversation turns, as shown in Figure 3, which presents accumulated metrics and win rates. Results revealed a clear upward trend in all key engagement metrics—especially verbosity, engagingness, interestingness, and listening—demonstrating that ChatWise’s effects strengthen as dialogues progress, while baseline gradually degrades, which highlights ChatWise’s capability in enhancing long-term conversational engagement.
(a) Verbosity.
(b) Engagingness Win Rate.
(c) Interestingness Win Rate.
(d) Listening Win Rate.
Figure 3: ChatWise performance with o3-mini as strategy provider. W/o ChatWise is the baseline and GPT-4o servers as utterance generator in all settings. Evaluation starts by skipping two warming-up turns (Sec 3).
4.5 Ablation Study
We evaluated ChatWise without user emotion in the strategy provider’s output (Table 2). Its consistent performance confirms that strategies are the primary driver of engaging conversations.
Strategy Provider GPT-4o GPT-4o w/o emotion w/o ChatWise
Verbosity 0.8635 0.8463 0.7398
Table 2: User verbosity under different strategy providers. GPT-4o w/o emotion keeps the strategy provider with user emotion information removed.
4.6 ChatWise’s Robustness to User Persona
Figure 4 shows the significant gain of ChatWise interacting with all simulated users, indicating its robustness to support varying senior characteristics.
Figure 4: Users’ dialogue verbosity (log-normalized, Appendix A.5) across 9 digital twins using different LLMs as strategy providers, compared with baseline (w/o ChatWise).
4.7 Dialogue Analysis
We identified 4 digital twins for which ChatWise was most effective and collected 40 additional dialogues from each to conduct a deeper analysis, with the main findings summarized below.
Transient user emotions: We defined emotion transition triplets as $(\text{e}_{i},\text{s},\text{e}_{i+1})$, where $\text{e}_{i}$ ($\text{e}_{i+1}$) represents the user’s emotion at turn $i$ ($i+1$), and $s$ is the strategy provided by ChatWise at turn $i$. We computed the average occurrence of emotion triplets and showed the top 15 most frequent in Figure 5. Most triplets reflect unchanged user emotions, indicating the challenge of detecting or influencing emotional shifts within a short turn.
Long-term user emotions: We analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 6). Over 48% users experienced emotional shifts post-dialogue, with significantly more positive changes after engaging with ChatWise. This implies that while transient emotions are difficult to track, strategic conversational support can improve user emotions over time.
Predominant strategies: We identified the top 10 most frequent strategies across digital twins and found that predominant interaction strategies remained consistent across user types (See Appendix A.8). Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations, suggesting their potential efficacy in fostering engaging dialogues.
Figure 5: Average occurrence of each emotion transition triplet across the samples of each digital twin.
Figure 6: Occurrence of user emotional changes. For instance, “neutral-joy” indicates user emotion transitioning from neutral (conversation begins) to joy (ends).
5 Conclusion and Future Work
We introduce ChatWise, a tuning-free yet principled framework for developing AI chatbots that foster long-term user engagement. Our dual-level generation design integrates clinical insights into LLM reasoning and is compatible with various LLMs. Extensive evaluations based on de-identified clinical data showed that our method robustly enhances the engagement and cognitive status of seniors. Moving forward, we aim to 1) enhance multi-turn interaction via RL and 2) validate its effectiveness through real user studies.
Limitations
All digital twins were provided as fine-tuned GPT-3.5 APIs, and its high economic cost, at the time when this project was going, constrained the volume of dialogues we generated. To mitigate this, future work will explore training digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality. Additionally, our study relies on simulated dialogues rather than real user interactions. A real user study is needed to validate the system’s effectiveness in dynamic settings, which we plan to incorporate into our future research.
Appendix A Appendix
A.1 Prompt design
The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-D5E0, including:
  • Strategy provider system prompt.
  • Moderator initial system prompt.
  • Moderator system prompt with strategies.
  • Strategy provider system prompt for ablation study.
  • Moderator system prompt with strategies for ablation study.
  • Structured output class for OpenAI models as strategy provider.
  • Structured output class for OpenAI models as strategy provider in ablation study.
  • System prompt for GPT-4o to extract the strategy given by Llama3.1.
A.2 Data Generation Configuration
We used GPT-4o as the utterance generator in ChatWise and tested different LLM backbones as strategy providers. Considering the Llama3.1-405B does not support structured output, we applied GPT-4o as the strategy extractor to structure the strategy output of Llama3.1-405B. The following are the configurations of each model:
Utterance generator (GPT-4o):
n=1
max_tokens=1024
top_p=1
temperature=1
GPT-4o, o3-mini as strategy provider:
n=1
max_tokens=1024
top_p=1
temperature=1
response_format=Strategy
Llama3.1-405B as strategy provider:
top_p: 0.9
max_tokens: 1024
temperature: 0.6
presence_penalty: 0
frequency_penalty: 0
GPT-4o as strategy extractor:
n=1
max_tokens=1024
top_p=1
temperature=1
A.3 Win rate for w/o ChatWise
The Win rate for w/o ChatWise is defined as:
$\begin{split}1-\text{average}(&\text{WR}_{GPT-4o}\\ &+\text{WR}_{o3-mini}\\ &+\text{WR}_{Llama3.1-405B})\end{split}$
,where $\text{WR}_{X}$ is ChatWise’s Win rate against the baseline with $X$ as strategy provider.
A.4 Dialogue Acts
The map of strategy to its corresponding abbreviated tag is listed in Table3.
Strategy Tag
Acknowledge (Backchannel) Ack
Statement-non-opinion StaNo
Statement-opinion Sta
Affirmation and Reassurance Agr
Appreciation App
Conventional-closing ConC
Hedge H
Other Oth
Quotation Quo
Action-directive AcD
Collaborative Completion CoC
Restatement or Paraphrasing Rep
Offers Options Commits Off
Self-talk Sel
Apology Apo
Reflection of Feelings RoF
Yes-No-Question YNQ
Wh-Question WhQ
Declarative Yes-No-Question DYNQ
Open-Question OpQ
Or-Clause OrC
Conventional-opening CoO
Self-disclosure Sd
Providing Suggestions PS
Information I
Table 3: Dialogue Act to its corresponding tag.
There are two kinds of strategies: backward-looking and forward-looking. Backward-looking strategies reflect how the current utterance relates to the previous discourse. Forward-looking strategies reflect the current utterance constrains the future beliefs and actions of the participants and affects the discourse. Table 4 and Table 5 provide definitions and examples of each.
Strategy Definiton Example
StaNo A factual statement or descriptive utterance that does not include an opinion. Me, I’m in the legal department.
Ack A brief utterance that signals understanding, agreement, or active listening. Uh-huh.
Sta A statement that conveys a personal belief, judgment, or opinion. I think it’s great
Agr Affirm the help seeker’s strengths, motivation, and capabilities and provide reassurance and encouragement. That’s exactly it.
App An expression of gratitude, admiration, or acknowledgment of another’s effort or input. I can imagine.
ConC A formal or socially standard utterance signaling the end of a conversation. Well, it’s been nice talking to you.
H An expression that introduces uncertainty or qualification to a statement, often to soften its impact. I don’t know if I’m making any sense or not.
Oth Exchange pleasantries and use other support strategies that do not fall into the above categories. Well give me a break, you know.
Quo A direct or indirect repetition of someone else’s words. Albert Einstein once said, “Imagination is more important than knowledge.”
AcD A command, request, or suggestion directing someone to take action. Why don’t you go first
CoC A continuation or completion of someone else’s utterance in a collaborative manner. If we want to make it to the top of the mountain before sunset, we should…
Rep A simple, more concise rephrasing of the help-seeker’s statements that could help them see their situation more clearly. It sounds like you’re saying that you’re struggling to stay on top of your work, and it’s leaving you feeling overwhelmed.
Off A statement proposing choices, making a commitment, or offering to do something. I’ll have to check that out
Sel An utterance directed at oneself, often reflecting internal thought processes or problem-solving. What’s the word I’m looking for
Apo An expression of regret or asking for forgiveness. I’m sorry.
RoF Articulate and describe the help-seeker’s feelings. It sounds like you’re feeling really frustrated and drained because your efforts don’t seem to be paying off.
Table 4: Backward-looking strategies, definition, and example.
Strategy Definiton Example
YNQ A question expecting a binary (yes/no) response. Do you have to have any special training?
WhQ A question beginning with a wh-word (e.g., what, who, where), seeking specific information. Well, how old are you?
DYNQ A statement posed as a question, expecting a yes/no answer. So you can afford to get a house?
OpQ A broad question inviting a wide range of responses, often conversational. How about you?
OrC A question offering explicit alternatives, often in the form of “or.” or is it more of a company?
CoO A socially standard utterance used to initiate a conversation. How are you?
Sd Divulge similar experiences that you have had or emotions that you share with the help-seeker to express your empathy. I completely understand how you feel. I remember feeling the same way before my first big presentation at work. I was so anxious, but I found that practicing a few extra times really helped calm my nerves.
PS Provide suggestions about how to change, but be careful to not overstep and tell them what to do. You can keep a note to stop your idea from going.
I Provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions. Taking silver line from Washington D.C. to Dulles Intel Airport costs about 1 hour.
Table 5: Forward-looking strategies, definition, and example.
A.5 Log-normalization
The following is the log-normalization function, where y is the normalized result, x is the input variable.
$y=\ln{\left(4x+1\right)}$
A.6 Additional results for Accumulated Multi-Turn Analysis
ChatWise’s accumulated performance on Fluency, and Making Sense is shown in Figure 7.
(a) Fluency Across Turns
(b) Making sense Across Turns
Figure 7: ChatWise performance with o3-mini as strategy provider across turns. W/o ChatWise is the baseline that ChatWise is not applied. GPT-4o servers as moderator in all settings.
A.7 Primary Strategies
We calculated the average occurrence of each strategy across each digital twin and listed their top 10 most frequently occurring strategies, as shown in Figure 8.
Figure 8: Strategy occurrence across digital twins.
A.8 Personalized Dialogue Analysis
This section shows the personalized dialogue analysis. The Strategy occurrence across digital twins is shown in Figure 9. The Occurrence of each emotion transition triplet across digital twins is shown in Figure 10. The Open Question, Statement-non-opinion, and Acknowledgment strategies still dominate ChatWise driven conversations, suggesting their potential effectiveness in fostering engagement in conversations. The user’s emotion is detected as unchanged in most triplets, indicating the difficulty of altering or measuring the user’s emotional movement within a short turn.
(a) Digital Twin 3.
(b) Digital Twin 5.
(c) Digital Twin 6.
(d) Digital Twin 9.
Figure 9: Strategy occurrence across digital twins.
(a) Digital Twin 3.
(b) Digital Twin 5.
(c) Digital Twin 6.
(d) Digital Twin 9.
Figure 10: Occurrence of each emotion transition triplet across digital twins.